OpenStudy (anonymous):

Prove that if \(R\) is an integral domain and \(x^2=1\) for some \(x\in R\) then \(x=\pm1\).

OpenStudy (kinggeorge):

Well, if \(x^2=1\), then \(x^2+(-1)=0\) which factors as \((x+1)(x+(-1))=0\)Since an integral domain has no zero divisors, either \(x+1=0\implies x=-1\) or \(x+(-1)=0\implies x=1\).

OpenStudy (anonymous):

That works. Thanks George.

OpenStudy (kinggeorge):

You're welcome.